# Convolution and Cross-Correlation of Count Sketches Enables Fast Cardinality Estimation of Multi-Join Queries

### v1

@@ Abstract.
cardinality estimation; sketching; synopsis data structures
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Santiago, Chileprice: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Query optimizationccs: Theory of computation Streaming modelsccs: Theory of computation Sketching and samplingAccepted at the International Conference on Management of Data 2024
@@ 3.1. Key insight for preserving information
${\bm{\Pi}}^{(1)}_{\cdot,a}=$${\bm{\Pi}}^{(2)}_{\cdot,b}=$00-100$\circ$000+10$=$0000000-100***000+10$=$-10000pt$(2+3)\bmod{5}=0$
@@ 3.2. General formulation by example
(5) $X$ $=\sum_{j\in J}\prod_{k=0}^{r-1}{c}_{k,G_{k}(j)}$
(6) $G_{k}(j)$ =($=\left\lparen\sum_{u\in\Omega(R_{k})}j_{\Psi(u)}\right\rparen\bmod{m}$
@@
Computing the estimates naively using Equation 5 has an exponential time complexity with the number of graph components. However, this can be improved significantly by factorizing the problem. We can then rely on the fact that circular cross-correlation can be computed efficiently in $O(m\log m)$ time using the fast Fourier transform (FFT). To perform this efficient estimation process methodically, one starts by selecting any joined attribute, say attribute 4 in our example. The process now aggregates all the sketches towards attribute 4 to obtain the estimate. This is implemented as a depth first traversal of the join graph with attribute 4 as the root node of a rooted tree and attributes 0 and 3 as the leaves. The general procedure for combining sketches to efficiently compute an estimate of the query is provided in Algorithm 3. Figure 5 exemplifies the estimation process, where the circles denote the sketches of the relations and the vertical rectangle represents the Hadamard product identity. The solid and dashed arrows indicate the Hadamard product and circular cross-correlation, respectively. In the pseudocode, we ensure that the recursion only moves away from any selected root attribute $o\in[w]$ by keeping track of the visited nodes $V$. The functions (I)FFT denote the (inverse) fast Fourier transform. This procedure reduces the inference time complexity to $O(rm\log{m})$, that is, nearly linear with respect to the sketch size.
@@
${\bm{c}}_{0}$${\bm{c}}_{2}$${\bm{c}}_{1}$${\bm{c}}_{3}$
Figure 5. The estimation process of the example query in Figure 3 with attribute 4 as the root.
@@ Appendix A Mean and variance
$\operatorname{\mathbb{E}}\left[\prod_{k=0}^{r-1}\prod_{u\in\Omega(R_{k})}\prod_{v\in\Gamma(u)}s_{j,\{u,v\}}\lparen i_{u}\rparen s_{j,\{u,v\}}\lparen i^{\prime}_{u}\rparen\right]$
@@
$\operatorname{\mathbb{E}}[X]={}$ $\sum_{j\in J}\sum_{i\in I}{\mathcal{F}}_{0}(i)\cdots{\mathcal{F}}_{r-1}(i)$
$\operatorname{\mathbb{E}}\left[\prod_{k=0}^{r-1}\operatorname{\mathbf{1}}\left\lparen H_{k}(i)=G_{k}(j)\right\rparen\prod_{u\in\Omega(R_{k})}\prod_{v\in\Gamma(u)}s_{\{u,v\}}(i_{u})\right]$
@@
$\operatorname{\mathbb{E}}[X^{2}]={}$ $\sum_{i\in I}\sum_{i^{\prime}\in I}\left\lparen\prod_{k=0}^{r-1}{\mathcal{F}}_{k}(i){\mathcal{F}}_{k}(i^{\prime})\right\rparen$ )
@@
$\prod_{\{u,v\}\in E}\operatorname{\mathbf{1}}\left\lparen i_{u}=i_{v}\land i^{\prime}_{u}=i^{\prime}_{v}\right\rparen+\operatorname{\mathbf{1}}\left\lparen i_{u}=i^{\prime}_{u}\neq i_{v}=i^{\prime}_{v}\right\rparen$ )
$\qquad\quad+\operatorname{\mathbf{1}}\left\lparen i_{u}=i^{\prime}_{v}\neq i_{v}=i^{\prime}_{u}\right\rparen$ )

### v3

@@ Abstract.
Cardinality Estimation, Sketching, Synopsis Data Structures
copyright: rightsretainedjournal: PACMMODjournalyear: 2024journalvolume: 2journalnumber: 3 (SIGMOD)article: 129publicationmonth: 6doi: 10.1145/3654932ccs: Information systems Query optimizationccs: Theory of computation Streaming modelsccs: Theory of computation Sketching and sampling
@@ 2. Background
$X,\operatorname{\mathbb{E}}[X]$ Estimate and expected value
$\epsilon,\delta$ Error bound and confidence
@@ 3.1. Key insight for preserving information
${\bm{\Pi}}^{(1)}_{\cdot,a}=$${\bm{\Pi}}^{(2)}_{\cdot,b}=$00-100$\circ$000+10$=$0000000-100*000+10$=$-10000$(2+3)\bmod{5}=0$
@@ 3.2. General formulation by example
(5) $X=\sum_{j\in J}\prod_{k=0}^{r-1}{c}_{k,G_{k}(j)},\quad\text{with}\quad G_{k}(j)=\left\lparen\sum_{u\in\Omega(R_{k})}j_{\Psi(u)}\right\rparen\bmod{m}$
@@
Computing the estimates naively using Equation 5 has an exponential time complexity with the number of graph components. However, this can be improved significantly by factorizing the problem. We can then rely on the fact that circular cross-correlation can be computed efficiently in $O(m\log m)$ time using the fast Fourier transform (FFT). To perform this efficient estimation process methodically, one starts by selecting any joined attribute, say attribute 4 in our example. The process now aggregates all the sketches towards attribute 4 to obtain the estimate. This is implemented as a depth first traversal of the join graph with attribute 4 as the root node of a rooted tree and attributes 0 and 3 as the leaves. The general procedure for combining sketches to efficiently compute an estimate of the query is provided in Algorithm 3. In the pseudocode, we ensure that the recursion only moves away from any selected root attribute $o\in[w]$ by keeping track of the visited nodes $V$. The functions (I)FFT denote the (inverse) fast Fourier transform. This procedure reduces the inference time complexity to $O(rm\log{m})$, that is, nearly linear with respect to the sketch size.
@@ Appendix A Mean and variance
$\operatorname{Var}(X)\leq\frac{1}{m}\sum_{i\in I}\sum_{i^{\prime}\in I}{\mathcal{F}}_{0}(i)\cdots{\mathcal{F}}_{r-1}(i){\mathcal{F}}_{0}(i^{\prime})\cdots{\mathcal{F}}_{r-1}(i^{\prime})\operatorname{\mathbb{E}}\left[\prod_{k=0}^{r-1}\prod_{u\in\Omega(R_{k})}\prod_{v\in\Gamma(u)}s_{j,\{u,v\}}\lparen i_{u}\rparen s_{j,\{u,v\}}\lparen i^{\prime}_{u}\rparen\right]$
@@
$\operatorname{\mathbb{E}}[X]=\sum_{j\in J}\sum_{i\in I}{\mathcal{F}}_{0}(i)\cdots{\mathcal{F}}_{r-1}(i)\operatorname{\mathbb{E}}\left[\prod_{k=0}^{r-1}\operatorname{\mathbf{1}}\left\lparen H_{k}(i)=G_{k}(j)\right\rparen\prod_{u\in\Omega(R_{k})}\prod_{v\in\Gamma(u)}s_{\{u,v\}}(i_{u})\right]$
@@
[…]